# Cluster Catch Digraphs with the Nearest Neighbor Distance

### v1

@@ 5 Monte Carlo Experiments for CCD-based Clustering
By optimizing $\delta$ for each simulation setting, KS-CCDs consistently deliver superior performance, yielding ARI values and success rates (%) close to 1 across most simulation settings, resulting in high average silhouette indices (consistently above 0.55). However, in certain cases when $n=50$, KS-CCDs exhibit reduced efficacy when dealing with 5 clusters. For example, when $d=2$ and $n=50$, the success rate (%) declines from 0.998 to 0.622 as the number of clusters increases from 2 to 5, This decrease can be attributed to a drop in the intensities of individual clusters. Furthermore, the small proximity of the $4^{th}$ and $5^{th}$ clusters to the existing clusters increases the likelihood of these clusters being erroneously confused as one. Fortunately, when $n\geq 100$, KS-CCDs exhibit substantial improvement in handling 5-cluster data sets.
@@
Although RK-CCDs are not as effective as KS-CCDs across lower-dimensional settings ($d\leq 5$), they still maintain ARI values and success rates above 0.9 in most scenarios, indicating solid performance in low-dimensional space. However, similar to the behavior of KS-CCDs, RK-CCDs also suffer performance degradation in certain cases when there are 5 clusters. For instance, when $d=2$ and $n=50$, the success rate (%) drops drastically from 0.998 to 0.082 as the number of clusters increases from 2 to 5, attributable to the low intensities of individual clusters. Additionally, The effectiveness of RK-CCDs declines across high-dimensional space. When $d=10$, the success rates (%) decrease substantially, reaching a minimum of 0.494 for 5-cluster data sets when $n=500$. This deterioration is further exacerbated when $d=20$, with ARIs falling below 0.8 and success rates (%) smaller than 0.4 in most settings.
@@
Under low-dimensional space ($d\leq 5$), UN-CCDs demonstrate comparable performance to RK-CCDs, achieving ARIs and success rates exceeding 0.95 in most settings. Furthermore, UN-CCDs exhibit substantial improvement over RK-CCDs in high-dimensional space. When $d=10$, both ARIs and success rates approach 1, much higher than those achieved by RK-CCDs. The performance gap between RK-CCDs and UN-CCDs becomes even more pronounced when $d=20$. For example, with 3-cluster data sets where $n=200$, UN-CCDs yield performance metrics of 0.921 (ARI), 0.558 (Sil), and 0.802 (success rate), respectively. These values are substantially higher than the corresponding 0.597, 0.376, and 0.312 obtained by RK-CCDs. This improvement is attributed to the newly developed SR-MCT with NND.
@@
Assuming CSR, UN-CCDs and RK-CCDs perform well with uniform clusters. To evaluate the performance of CCD-based clustering methods on non-uniform data, we conduct simulations with Gaussian clusters. Each Gaussian cluster has a covariance matrix $\Delta*I_{d}$, where $\Delta$ is a random uniform variable between 0.8 and 1.2, and $I_{d}$ represents a $dxd$ identity matrix. All other settings, including dimensionality, data size, cluster centers, and the number of clusters, remain the same. Three realizations with different number of Gaussian clusters ($d=2$) are presented in Figure 3. The simulation results are summarized in Table 5, 7, and 8.
@@
Table 5: The simulation results of KS-CCDs on a set of synthetic data sets consisting of Gaussian clusters.
Density Parameter: $d]{\delta}$ $d=2$ $d=3$ $d=5$ $d=10$ $d=20$
$n=50$ 2 clusters 1.90 1.10 0.60 0.32 0.20
3 clusters 1.60 0.93 0.52 0.30 0.19
5 clusters 1.30 0.79 0.46 0.28 0.18
$n=100$ 2 clusters 3.00 1.40 0.71 0.35 0.21
3 clusters 2.30 1.20 0.65 0.32 0.20
5 clusters 1.90 0.99 0.58 0.32 0.17
$n=200$ 2 clusters 4.20 1.70 0.76 0.39 0.23
3 clusters 3.50 1.50 0.71 0.37 0.21
5 clusters 2.70 1.30 0.66 0.35 0.17
$n=500$ 2 clusters 7.1 2.5 0.95 0.43 0.24
3 clusters 5.6 2.1 0.90 0.41 0.24
5 clusters 4.3 1.8 0.82 0.41 0.19
Table 6: The optimal scaled densities ($d]{\delta}$) of KS-CCDs on synthetic data sets consisting of Gaussian clusters.
@@
Table 7: The simulation results of RK-CCDs on a set of synthetic data sets consisting of Gaussian clusters.
@@
Table 8: The simulation results of UN-CCDs on a set of synthetic data sets consisting of Gaussian clusters.
@@
While the optimal density parameter $\delta$ is selected for each simulation setting, KS-CCDs demonstrate reduced effectiveness with Gaussian clusters. In low-dimensional settings ($d\leq 5$), KS-CCDs deliver slightly lower performance compared to the scenarios with uniform clusters. Notably, ARI values and success rates remain high ($>0.95$), indicating high efficacy even with Gaussian clusters. However, KS-CCDs degrade substantially in high-dimensional settings ($d\geq 10$), particularly with three or more clusters. For example, when $d=20$ and $n=200$, KS-CCDs deliver average performance metrics of 0.736 (ARI), 0.257 (Sil), and 0.550 (success rate) for 3-cluster data sets, drastically lower than the corresponding 1, 0.587, and 1 obtained with uniform clusters. This performance decline can be explained by the decreased intensities of Gaussian clusters in high-dimensional space, an effect amplified with a larger number of clusters (e.g., 5 clusters). Moreover, Gaussian clusters are unbounded, further contributes to the performance reduction.
@@
Similar to the simulations with uniform clusters, RK-CCDs generally perform well in low-dimensional space, with ARIs and success rates exceeding 0.9 except certain cases with 5 clusters, where the proximities between clusters and the intensities of individual clusters are small. For instance, with $d=2$, $n=50$, and 5 clusters, the performance metrics of RK-CCDs are 0.627 (ARI), 0.535 (Sil), and 0.234 (success rate), respectively. RK-CCDs demonstrate similar behavior in high-dimensional space ($d\geq 10$). Their performance is slightly worse than KS-CCDs when $d=10$, and deteriorate further when $d=20$, increasing the data size does not help. For example, when $d=20$ and $n=500$, the ARI and success rate are 0.297 and 0, respectively, due to the limitations of the SR-MCT based on the Ripley’s K function.
@@
Similar to the previous analysis, UN-CCDs perform similarly to RK-CCDs and KS-CCDs in low-dimensional space ($d\leq 5$), Both ARIs and success rates exceed 0.9 in most scenarios, with a few exceptions observed when dealing with 5 clusters. These results are slightly lower than those obtained with uniform clusters. However, UN-CCDs yield superior overall performance in high-dimensional space ($d\geq 10$). When $d=10$, UN-CCDs deliver comparable performance to KS-CCDs, outperforming RK-CCDs, as evidenced by higher ARIs and success rates. To illustrate, with $n=200$, $d=10$, and 3 clusters, the ARIs for KS-CCDs, RK-CCDs, and UN-CCDs are 0.963, 0.877, and 0.955, respectively. Furthermore, UN-CCDs exhibit a substantial advantage over the other two CCDs when $d=20$. For example, with $n=200$, $d=20$, and 5 clusters, the ARIs for KS-CCDs, RK-CCDs, and UN-CCDs are 0.263, 0.342, and 0.439, respectively.
@@
To summarize, KS-CCDs demonstrate the best overall performance with uniform clusters when the optimal density parameter $\delta$ is selected, achieving ARI and success rates close to 1 in most scenarios, invariant to the dimensionality $d$. UN-CCDs exhibit slightly lower performance compared to KS-CCDs, and show a substantial advantage over RK-CCDs when $d\geq 10$. All three CCDs perform slightly worse with Gaussian clusters in low-dimensional space ($d\leq 5$), further deteriorate when $d\geq 10$, where UN-CCDs outperform both RK-CCDs and KS-CCDs. The performance of all three CCDs declines with 5 clusters, particularly when the data size $n$ is small. This is attributed to the low intensities of individual clusters and the small proximities between clusters.
6 Real and Complex Data sets
@@
We assess the performance of three CCD-based clustering methods using real and complex artificial data sets in this section, and compare them with some well-known, established methods. Moreover, we want to explore if UN-CCDs outperform RK-CCDs after multiple enhancements. The data sets, obtained from UCI Machine Learning Repository and Clustering basic benchmark [16, 22], are generally more complex and challenging than the data sets used in Sections 5. To prevent features with larger values from dominating, we preprocess the data sets by normalizing them (except the synthetic data sets including asymmetric, R15, and D31).
The details of each data set are summarized below. Moreover, we present the asymmetric, R15, and D31 data sets in Figure 4.
@@
  • asymmetric: A synthetic 2-dimensional data sets with 5 elliptical Gaussian clusters of unbalanced cluster size and intensities, and the inter-cluster distances varies [47].
  • R15 and D31: Synthetic data sets consist of 15 or 31 similar Gaussian clusters [55]
$n$ $d$ of clusters
@@
We compare the performance of CCDs with established clustering algorithms, Density Based Spatial Clustering of Applications with Noise (DBSCAN) [19], the Minimal Spanning Tree (MST) Method [63], Spectral Clustering (Spectral) [40], K-means++ [3], and Louvain Clustering Method (Louvain) [8].
@@
DBSCAN is a density-based clustering method proposed by Ester et al. [19], often used to handle data sets with noise or outliers. It identifies clusters by finding “seeds”, which are the points deep inside clusters that have a minimum number (denoted as MinPts) of neighbors within a given radius (denoted as Eps). After that, it constructs clusters by finding all the points that are density-reachable from seeds. Finally, the points not connected to any seeds are labeled as noise or outliers. MinPts and Eps are the two input parameters, we follow a heuristic offered by Ester et al., which sets MinPts to 4, and find the value at the first “elbow” of the sorted $4$-dist (the distance of a point to its $4^{th}$ nearest neighbor) plot, setting it to Eps [19].
@@
The MST clustering method employs a graph-based approach [63]. This method involves constructing a graph by connecting all data points with the minimum sum of edge weights, typically represented by the distance between two points. Subsequently, edges with significantly larger weights than their adjacent edges’ average weight are identified as ”inconsistent” and removed. This removal effectively breaks the MST into subtrees, each representing a distinct cluster. We set the threshold value for “inconsistent” edges to 2, which delivers better overall performance compared to other thresholds.
Spectral clustering is a powerful method, employing eigenvalue decomposition of the Laplacian matrix, which are derived from the similarity matrix of the data set [40]. Finally, it applies a traditional method for clustering, and we choose $K$-means in this experiment, with the number of clusters known.
@@
$K$-means++ enhances the classic $k$-means clustering method [3]. Sensitivity to the initial chosen cluster centroids is a key weakness of $k$-means, $k$-means++ address this limitation by selecting initial centroids that are far apart leading to better and stable clustering results and faster convergence. Similar to spectral clustering, we conduct $k$-means++ with the number of clusters known.
@@
Louvain clustering method is a greedy algorithm finding communities in large networks [8]. It merges communities or notes recursively by maximizing the “modularity”, which is a measure of the quality of a partition. The algorithm stops until no further improvement in modularity can be found. Although having a optional resolution parameter affects the size and number of communities, the basic Louvain clustering method does not require any input parameters. In this experiment, we employ the standardized Gaussian kernel to construct a network for the data set and apply Louvain clustering method with the default resolution parameter (i.e., 1).
@@
Consider the CCD-based methods, we tune the significance level $\alpha$ of SR-MCT for RK-CCDs and UN-CCDs. Specifically, $\alpha$ is set to $1\%$ for both methods For KS-CCDs, we choose the $d]{\delta}$ value that maximizes the average silhouette index in each data set.
@@
We choose Adjust Rand index and Silhouette Index as external and internal validation measures. Furthermore, we record the number of clusters detected by each method. The performance of all methods are summarized in Table 10.
iris ($k$=3) seeds ($k$=3) knowledge ($k$=3) wholesale ($k$=3) asymmetric ($k$=5) R15 ($k=15$) D31 ($k=31$)
@@
All the methods obtain comparable ARIs on the iris data set. Precisely, Louvain and UN-CCDs attain the highest ARIs of 0.630 and 0.599, respectively. KS-CCDs and DBSCAN produce the best Sils. RK-CCDs, UN-CCDs, and Louvain clustering method correctly identify the true number of clusters.
@@
In the seeds data set, K-means++, spectral clustering, and Louvain clustering method perform better than other methods when considering ARI. Meanwhile, UN-CCDs and KS-CCDs achieve the highest Sils of 0.468 and 0.467, respectively, indicating better-defined cluster structures. Both RK-CCDs and Louvain clustering method successfully identify the correct number of clusters.
@@
In the knowledge data set, The performance of all the methods is generally mediocre. K-mean++ and spectral clustering get the highest ARIs, 0.202 and 0.189, respectively. KS-CCDs and DBSCAN produce the best Sils. MST is the only method that identified 3 clusters.
@@
All the methods perform poorly on the wholesale data set, Suggested by the ARIs, which are around 0, indicating the clustering results akin to random assignment. Therefore, we turn to Sil for evaluation, KS-CCDs and DBSCAN outperform other methods, achieving Sil of 0.765 and 0.569, respectively. Despite having the second-lowest Sil score, UN-CCDs is the only method that correctly identifies the number of clusters.
@@
When the number of clusters is known, K-means++ and spectral clustering exhibit the best performance. These methods achieve ARI values of 0.882 and 0.975, respectively, and Sil values of 0.645 and 0.623, respectively. The CCD-based methods show slightly lower performance, with ARI and Sil around 0.7 and 0.6, respectively. However, these methods identify 4 clusters, the closest to the actual number.
@@
The three CCD-based algorithms perform better than the other methods on the R15 data set, they successfully capture 15 Gaussian clusters, achieving ARIs close to 1 and Sils around 0.75. DBSCAN, K-means++, and spectral clustering deliver slightly inferior performance, with ARIs around 0.9 and Sils around 0.65.

### v2

@@ 4.1 Experiment with Uniform Settings
According to the simulation result, KS-CCDs consistently deliver superior performance when the optimal $\delta$ is selected, yielding ARI values and success rates (SR) close to 1 across most simulation settings, resulting in high average Sil values (consistently above 0.55). However, in certain cases when $n=50$, KS-CCDs exhibit reduced efficiency when dealing with 5 clusters. For example, when $d=2$ and $n=50$, the success rate (SR) declines from 0.998 to 0.622 as the number of clusters increases from 2 to 5, As shown in the Figure 3, the ARI value also decreases when the number of cluster increase. The degradation in performance can be attributed to a drop in the intensities of individual clusters. Furthermore, the small proximity of the $4^{th}$ and $5^{th}$ clusters to the existing clusters increases the likelihood of these clusters being erroneously confused as one. Fortunately, when $n\geq 100$, KS-CCDs exhibit substantial improvement in handling 5-cluster datasets.
@@
Although RK-CCDs are not as effective as KS-CCDs across lower-dimensional settings ($d\leq 5$), they still maintain ARI values and success rates (SR) above 0.9 in most scenarios, indicating solid performance in low-dimensional space. However, similar to the behavior of KS-CCDs, RK-CCDs also suffer performance degradation in certain cases when there are 5 clusters. For instance, when $d=2$ and $n=50$, the success rate (SR) drops drastically from 0.998 to 0.082 as the number of clusters increases from 2 to 5, attributable to the low intensities of individual clusters. Additionally, RK-CCDs perform worse in high-dimensional space, due to the limitations of the Ripley’s $K$ function when $d$ is large. for example, when $d=10$, the success rates (SR) decrease substantially, reaching a minimum of 0.494 for 5-cluster datasets when $n=500$. This deterioration is further exacerbated when $d=20$, with ARI values falling below 0.8 and success rates (SR) smaller than 0.4 in most settings.
@@
Under low-dimensional space ($d\leq 5$), UN-CCDs demonstrate comparable performance to RK-CCDs, achieving ARI values and success rates (SR) exceeding 0.95 in most settings. Furthermore, UN-CCDs exhibit substantial improvement over RK-CCDs in high-dimensional space. When $d=10$, both ARI and success rate approach 1, much higher than those achieved by RK-CCDs. The performance gap between RK-CCDs and UN-CCDs becomes even more pronounced when $d=20$. For example, with 3-cluster datasets where $n=200$, UN-CCDs yield performance metrics of 0.921 (ARI), 0.558 (Sil), and 0.802 (success rate), respectively. These values are substantially higher than the corresponding 0.597, 0.376, and 0.312 obtained by RK-CCDs. As we talked previously in Section 3.2, this improvement is attributed to the newly developed MC-SRT, which employed NND instead of the Ripley’s $K$ function with several additional key enhancements.
4.2 Experiment with Gaussian Settings
@@
Assuming CSR, UN-CCDs and RK-CCDs perform well with uniform clusters. To evaluate the performance of CCD-based clustering methods on non-uniform data, we conduct simulations with Gaussian clusters. Each Gaussian cluster has a covariance matrix $\Delta*I_{d}$, where $\Delta$ is a random uniform variable between 0.8 and 1.2 (for random scaling), and $I_{d}$ represents a $d\times d$ identity matrix. All other settings, including dimensionality, data size, cluster centers, and the number of clusters, remain the same as in the uniform settings. Three realizations with different number of Gaussian clusters ($d=2$) are presented in Figure 6. The simulation results are summarized in Table 7, 8, and 9. Similarly, we present the performance trend (ARIs) of the CCD-based methods from Figures 7 to 9.
(a) 2 clusters
(b) 3 clusters
(c) 5 clusters
@@
Density Parameter: $\sqrt[d]{\delta}$ $d=2$ $d=3$ $d=5$ $d=10$ $d=20$
$n=50$ 2 clusters 1.90 1.10 0.60 0.32 0.20
3 clusters 1.60 0.93 0.52 0.30 0.19
5 clusters 1.30 0.79 0.46 0.28 0.18
$n=100$ 2 clusters 3.00 1.40 0.71 0.35 0.21
3 clusters 2.30 1.20 0.65 0.32 0.20
5 clusters 1.90 0.99 0.58 0.32 0.17
$n=200$ 2 clusters 4.20 1.70 0.76 0.39 0.23
3 clusters 3.50 1.50 0.71 0.37 0.21
5 clusters 2.70 1.30 0.66 0.35 0.17
$n=500$ 2 clusters 7.1 2.5 0.95 0.43 0.24
3 clusters 5.6 2.1 0.90 0.41 0.24
5 clusters 4.3 1.8 0.82 0.41 0.19
Table 6: The optimal scaled densities ($\sqrt[d]{\delta}$) of KS-CCDs on synthetic datasets consisting of Gaussian clusters.
@@
Table 7: The simulation results of KS-CCDs on a set of synthetic datasets consisting of Gaussian clusters. While highly efficient with Gaussian clusters in low dimensions, KS-CCDs’ performance degrades substantially in high-dimensional spaces due to the decreased intensity and unbounded nature of Gaussian distributions.
Figure 7: The line plots of the ARIs of KS-CCDs, under the Gaussian cluster settings.
@@
Table 8: The simulation results of RK-CCDs on a set of synthetic datasets consisting of Gaussian clusters. Similar to Uniform settings, the performance of RK-CCDs deteriorates significantly in high-dimensional space due to the limitations of the Ripley’s $K$ function.
Figure 8: The line plots of the ARIs of RK-CCDs, under the Gaussian cluster settings.
@@
Table 9: The simulation results of UN-CCDs on a set of synthetic datasets consisting of Gaussian clusters. The UN-CCD method demonstrate substantial advantage over KS-CCDs and RK-CCDs in high-dimensional space.
Figure 9: The line plots of the ARIs of UN-CCDs, under the Gaussian cluster settings.
@@
While the optimal density parameter $\delta$ is selected for each simulation setting, KS-CCDs demonstrate reduced effectiveness with Gaussian clusters. In low-dimensional settings ($d\leq 5$), KS-CCDs deliver slightly lower performance in the comparable scenarios with uniform clusters. Notably, ARI values and success rates (SR) remain high (>0.95>0.95), indicating high efficiency even with Gaussian clusters. However, as shown in the second line plot of Figure 7. KS-CCDs degrade substantially in high-dimensional settings ($d\geq 10$), particularly with three or more clusters. For example, when $d=20$ and $n=200$, KS-CCDs deliver average performance metrics of 0.736 (ARI), 0.257 (Sil), and 0.550 (success rate) for 3-cluster datasets, drastically lower than the corresponding 1, 0.587, and 1 obtained with uniform clusters. This performance decline can be explained by the decreased intensities of Gaussian clusters in high-dimensional space, an effect amplified with a larger number of clusters (e.g., 5 clusters), suggested by the third line plot of Figure 7. Moreover, the support of Gaussian clusters is unbounded, further contributes to the performance degradation.
@@
Similar to the simulations with uniform clusters, RK-CCDs generally perform well in low-dimensional space, with ARI values and success rates (SR) exceeding 0.9 except certain cases with 5 clusters, where the proximities between clusters and the intensities of individual clusters are small. For instance, with $d=2$, $n=50$, and 5 clusters, the performance metrics of RK-CCDs are 0.627 (ARI), 0.535 (Sil), and 0.234 (success rate), respectively. RK-CCDs demonstrate similar behavior in high-dimensional space ($d\geq 10$) as in the uniform settings. Their performance is slightly worse than KS-CCDs when $d=10$, and deteriorate further when $d=20$, increasing the data size does not help. For example, when $d=20$ and $n=500$, the ARI and success rate are 0.297 and 0, respectively, due to the limitations of the MC-SRT based on the Ripley’s K function.
@@
Similar to the previous analysis, UN-CCDs perform similarly to RK-CCDs and KS-CCDs in low-dimensional space ($d\leq 5$), Both ARI values and success rates exceed 0.9 in most scenarios, with a few exceptions observed when dealing with 5 clusters. These results are slightly lower than those obtained with uniform clusters. However, UN-CCDs yield superior overall performance in high-dimensional space ($d\geq 10$). When $d=10$, UN-CCDs deliver comparable performance to KS-CCDs, outperforming RK-CCDs, as evidenced by higher ARI values and success rates. To illustrate, with $n=200$, $d=10$, and 3 clusters, the ARI values for KS-CCDs, RK-CCDs, and UN-CCDs are 0.963, 0.877, and 0.955, respectively. Furthermore, UN-CCDs exhibit a substantial advantage over the other two CCDs when $d=20$. For example, with $n=200$, $d=20$, and 5 clusters, the ARI values for KS-CCDs, RK-CCDs, and UN-CCDs are 0.263, 0.342, and 0.439, respectively.
@@
To summarize, KS-CCDs demonstrate the best overall performance with uniform clusters when the optimal density parameter $\delta$ is selected, achieving ARI values and success rates close to 1 in most scenarios, invariant to the dimensionality $d$. UN-CCDs exhibit slightly lower performance compared to KS-CCDs, and show a substantial advantage over RK-CCDs when $d\geq 10$. All three CCDs perform slightly worse with Gaussian clusters in low-dimensional space ($d\leq 5$), further deteriorate when $d\geq 10$, where UN-CCDs outperform both RK-CCDs and KS-CCDs. The performance of all three CCDs declines with 5 clusters, particularly when the data size $n$ is small. This is attributed to the low intensities of individual clusters and the small proximities between clusters.
4.3 Experiment on Datasets with Noise
In this section, we conduct additional simulation experiments to evaluate whether the UN-CCD method is robust to noise. The general simulation setup is similar to the experiments with Gaussian settings, where each Gaussian cluster has a random scaling covariance matrix $\Delta*I_{d}$. To simplify the study, we set the number of clusters to 3, and the data size to 200 plus the number of noise points. Noise refers to the points that are randomly distributed within the scope of the entire dataset. we simulate datasets with varying noise level ($3\%$, $5\%$, $10\%$, $20\%$) to assess the performance trend of UN-CCD methods. Additionally, we consider both 3-dimensional and 10-dimensional datasets to study the impact of noise level with different dimensions. Specific details are outlined below. Some realizations of datasets with Gaussian clusters in 2-dimensional space (although the simulation experiments are conducted on 3 and 10-dimensional space) are presented in Figure 10 (for illustration purposes). Similar to the previous simulation study, Each simulation setting is repeated 1000 times to ensure accurate results. The simulation results are summarized in Table 10, similar to the previous simulations, we record the average ARI, Sil, and SR under each setting.
  • i.
    The dimensionality ($d$) of the simulated datasets: 3, 10;
  • ii.
    The noise levels ($l$): $3\%$, $5\%$, $10\%$, $20\%$
  • iii.
    The size of datasets ($n$): $200\times(1+l)$;
  • iv.
    The size of each cluster is equal;
  • v.
    The number of clusters: 3;
  • vi.
    The centers of Gaussian clusters: $\bm{\mu_{1}}=(\underbrace{3,...,3}_{d})$, $\bm{\mu_{2}}=(9,\underbrace{3,...,3}_{d-1})$, and $\bm{\mu_{3}}=(3,9,\underbrace{3,...,3}_{d-2})$.
  • vii.
    The covariance matrix of each Gaussian cluster: $\Delta*I_{d}$, where $\Delta\sim U[0.8,1.2]$ and $I_{d}$ is a $d$-dimensional identity matrix.
(a) $3\%$
(b) $5\%$
(c) $10\%$
(d) $20\%$
Figure 10: Some realizations of the simulation studying the impact of different noise levels, the noise level varies from $3\%$ to $20\%$. Red crosses are noise, black points are regular observations. The noise levels are indicated below each sub-figure.
UN-CCDs $l=3\%$ $l=5\%$ $l=10\%$ $l=20\%$
ARI Sil SR ARI Sil SR ARI Sil SR ARI Sil SR
$d=3$ 0.977 0.619 1.000 0.977 0.619 1.000 0.975 0.618 0.998 0.971 0.617 0.982
$d=10$ 0.959 0.386 0.973 0.961 0.384 0.977 0.954 0.384 0.960 0.947 0.385 0.942
Table 10: The simulation results of UN-CCDs on a set of synthetic datasets with varying noise level and 3 Gaussian clusters, demonstrating that the UN-CCD method is invariant to varying noise levels.
The UN-CCD method is highly robust to noise, maintaining consistent performance across various noise levels. Specifically, when $d=3$, increasing the noise level ($l$) from $3\%$ to $20\%$ has a negligible effect. For example, the ARI values are 0.977, 0.977, 0.975, and 0.971 when $l$ increases from $3\%$ to $20\%$; meanwhile, the Success Rate (SR) only slightly decreases from 1.000 to 0.982. While the effect of noise is greater when $d=10$, the method remains effective, with the SR decreasing from 0.973 to 0.942. Overall, these results demonstrate the UN-CCD’s reliable performance in the presence of noise across various dimensions.
5 Real Data Examples
@@
We assess the performance of three CCD-based clustering methods using some benchmark real and synthetic datasets in this section, and compare them with some well-known, established clustering methods. Moreover, we want to explore if UN-CCDs outperform RK-CCDs after multiple enhancements. The datasets, obtained from UCI Machine Learning Repository and Clustering basic benchmark [uci_repository, ClusteringDatasets], are generally more complex in the data structure, thus more challenging for clustering than the datasets used in Section 4. To prevent features with larger scales from dominating, we preprocess the datasets by normalizing them (except the synthetic datasets, including “asymmetric”, “R15”, and “D31” datasets).
The details and background of each dataset are summarized below. Moreover, we present the “asymmetric”, “R15”, and “D31” datasets in Figure 11. We are treating all datasets unlabeled, and compare the obtained clusters with the classes already available in the datasets.
@@
  • asymmetric: A synthetic 2-dimensional datasets with 5 elliptical Gaussian clusters of unbalanced cluster size and intensities, and the inter-cluster distances varies [rezaei2020can].
  • R15 and D31: Synthetic datasets consist of 15 or 31 similar Gaussian clusters [veenman2002maximum]
$n$ $d$ # of clusters
@@
We compare the performance of CCDs with established clustering algorithms, Density Based Spatial Clustering of Applications with Noise (DBSCAN) [ester1996density], the Minimal Spanning Tree (MST) Method [MST], Spectral Clustering (Spectral) [ng2001spectral], K-means++ [arthur2006k], and Louvain Clustering Method (Louvain) [blondel2008fast].
@@
DBSCAN is a density-based clustering method proposed by Ester et al. [ester1996density], often used to handle datasets with noise or outliers. It identifies clusters by finding “seeds”, which are the points deep inside clusters that have a minimum number (denoted as MinPts) of neighbors within a given radius (denoted as Eps). After that, it constructs clusters by finding all the points that are density-reachable from seeds. Finally, the points not connected to any seeds are labeled as noise or outliers. MinPts and Eps are the two input parameters, we follow a heuristic offered by Ester et al., which sets MinPts to 4, and find the value at the first “elbow” of the sorted $4$-dist (the distance of a point to its $4^{th}$ nearest neighbor) plot, setting it to Eps [ester1996density].
@@
The MST clustering method employs a graph-based approach [MST]. This method involves constructing a graph by connecting all data points with the minimum sum of edge weights, typically represented by the distance between two points. Subsequently, edges with significantly larger weights than their adjacent edges’ average weight are identified as ”inconsistent” and removed. This removal effectively breaks the MST into subtrees, each representing a distinct cluster. After the trial-and-error process, We set the threshold value for “inconsistent” edges to 2, which delivers the best overall performance compared to other thresholds.
Spectral clustering starts by constructing a similarity matrix of the given dataset [ng2001spectral]. Then, it derives the Laplacian matrix from the similarity matrix, and the eigenvalues of the Laplacian matrix provide a low-dimensional embedding of the original data. Finally, it applies a traditional method for clustering. In our experiment, we chose Gaussian similarity function to define the similarity and $K$-means for clustering, with the number of clusters $k$ as the input parameter.
@@
$K$-means++ enhances the classic $K$-means clustering method [arthur2006k]. A key weakness of $K$-means is the sensitivity to the initially chosen cluster centroids. $K$-means++ addresses this limitation by selecting initial centroids that are far apart, leading to better and more stable clustering results and faster convergence. Similar to spectral clustering, we conduct $K$-means++ on each dataset with the true number of clusters as an input.
@@
Louvain clustering method is a greedy algorithm finding communities in large networks by maximizing the “modularity”, which measures the quality of a partition[blondel2008fast]. The algorithm first optimizes modularity by iteratively moving individual nodes between communities, then aggregates these communities into new nodes and repeats the process until no further improvement is possible. In this experiment, we first employ the standardized Gaussian kernel to construct a network for the given dataset, and then apply the Louvain clustering method with the default resolution parameter to find the communities (i.e., 1).
@@
Consider the CCD-based methods, $\alpha$ is set to $1\%$ for all the dataset for simplicity. For KS-CCDs, we choose the $\sqrt[d]{\delta}$ value that maximizes the average silhouette index in each dataset.
@@
We choose the Adjust Rand Index and Silhouette Index as external and internal validation measures. Furthermore, we record the number of clusters detected by each method. The performance of all methods is summarized in Table 12.
iris (kk=3) seeds (kk=3) knowledge (kk=3) wholesale (kk=3) asymmetric (kk=5) R15 ($k=15$) D31 ($k=31$)
@@
Except MST, all the methods obtain comparable ARIs on the “iris” dataset. Precisely, Louvain and UN-CCDs attain the highest ARIs of 0.630 and 0.599, respectively. KS-CCDs and DBSCAN produce the best Sil values. RK-CCDs, UN-CCDs, and Louvain clustering method correctly identify the true number of clusters. MST’s underperformance is due to its sensitivity to the varying intensities of the iris clusters.
“seeds” datasets:
@@
In the “seeds” dataset, $K$-means++, spectral clustering, and Louvain clustering method perform better than other methods when considering ARI. On the other hand, UN-CCDs and KS-CCDs achieve the highest Sil values of 0.468 and 0.467, respectively, indicating better-defined cluster structures. Both RK-CCDs and Louvain clustering method successfully identify the correct number of clusters. DBSCAN underperforms due to the difficulty to find a single Eps value that works globally.
“knowledge” dataset:
@@
In the “knowledge” dataset, The performance of all the methods is generally mediocre. $K$-mean++ and spectral clustering get the highest ARI values, 0.202 and 0.189, respectively. KS-CCDs and DBSCAN produce the best Sil values. MST is the only method that identified 3 clusters.
“wholesale” dataset:
@@
All the methods perform poorly on the “wholesale” dataset, Suggested by the ARIs, which are around 0, indicating the clustering results akin to random assignment. Therefore, we consider Sil for evaluation, KS-CCDs and DBSCAN outperform other methods, achieving Sil values of 0.765 and 0.569, respectively. Despite having the second-lowest Sil value (0.113), UN-CCDs is the only method that correctly identifies the number of clusters.
“asymmetric” dataset:
@@
When the number of clusters is known, $K$-means++ and spectral clustering exhibit the best performance on the “asymmetric” dataset. These methods achieve ARI values of 0.882 and 0.975, respectively, and Sil values of 0.645 and 0.623, respectively. The CCD-based methods show slightly lower performance, with ARI and Sil values around 0.7 and 0.6, respectively. However, these methods identify 4 clusters, the closest to the actual number.
“R15” dataset:
@@
The three CCD-based methods perform better than the other methods on the “R15” dataset, they successfully capture 15 Gaussian clusters, achieving ARI values close to 1 and Sil values around 0.75. DBSCAN, $K$-means++, and spectral clustering deliver slightly inferior performance, with ARI values around 0.9 and Sil values around 0.65.
“D31” dataset: